OpenStudy (anonymous):

One circular table has a diameter of 9 ft, and another circular table has a diameter of 20 ft. How much greater is the area of the larger table? Round to the nearest whole number.

OpenStudy (davidusa):

What is the radius of each of the 2 circles? (divide diameter by 2)

OpenStudy (davidusa):

10 and 4.5

OpenStudy (anonymous):

about 251 ft2 about 146 ft2 about 336 ft2 about 1,002 ft2 theses are the answer choices

OpenStudy (davidusa):

So first to find area of circle, what do u do?

OpenStudy (anonymous):

What is the equation for area of a circle? Divide the diameter by 2 to find the radius. Plug into the equation. Subtract the 2 areas (1 for each circle)

OpenStudy (davidusa):

100pi-20.25pi

OpenStudy (davidusa):

use a calc for that.

OpenStudy (anonymous):

You need to use the following equations: Area of circle = 2*pi*(r^2) radius, r = diameter/2 Plug in the given values and subtract the smaller table's area from the larger table's area. It is about 251ft^2.

OpenStudy (wolf1728):

Area = 1/4 (PI * d²)

OpenStudy (davidusa):

not 1/4

OpenStudy (anonymous):

8. A team in science class placed a chalk mark on the side of a wheel and rolled the wheel in a straight line until the chalk mark returned to the same position. The team then measured the distance the wheel had rolled and found it to be 35 cm. To the nearest tenth, what is the area of the wheel? 27.5 cm2 97.5 cm2 195.1 cm2 390.1 cm2 this question i asked earlier

OpenStudy (anonymous):

thank youuu i tried itt .

OpenStudy (wolf1728):

9² = 81 20² = 400 PI * 81 = 254.4690049408 PI * 400 = 1,256.6370614359 Difference = 1002.1680564952

OpenStudy (davidusa):

NO. those are using the diameters. its radius squared.

OpenStudy (davidusa):

CLASSIC MISTAKE!!!

OpenStudy (wolf1728):

David Circle Area = PI * r² = 1/4 * (PI * d²)

OpenStudy (davidusa):

OH GOSH. just do it ez way ok?

OpenStudy (anonymous):

i got the first answer already what abt 8 im stuck /:

OpenStudy (anonymous):

about question 8 i put *

OpenStudy (wolf1728):

Okay - anyway the answer is 1,002 Okay now for question 8

OpenStudy (anonymous):

so the 1st questioni s 1,002 ft2 not 251 ft2

OpenStudy (wolf1728):

35 cm = wheel circumference wheel diameter = 35/PI = 11.1408460164 wheel area = 1/4 PI*d² area = 389.9296105751

OpenStudy (wolf1728):

Yes first answer is 1,002 - look at my work

OpenStudy (anonymous):

yea i see thank you

OpenStudy (wolf1728):

u r welcome and I think I answered question 8 properly too

OpenStudy (anonymous):

35 cm ?

OpenStudy (wolf1728):

area = 390

OpenStudy (wolf1728):

OR if you use PI = 22/7 then area = 390.086557023

OpenStudy (wolf1728):

Incidentally, I believe the record for computing PI is over a trillion digits!!! That's why I don't like using 22/7. Heck can't we at least keep it accurate to 3.14159? LOL

OpenStudy (anonymous):

lol are you a math teacher ?

OpenStudy (wolf1728):

No - just someone who was good at math when I was in school a long time ago.

OpenStudy (anonymous):

60 cm 30 cm 90 cm 120 cm what about this question ?

OpenStudy (anonymous):

The circumference of a circle is 60π cm. What is the radius of the circle?

OpenStudy (wolf1728):

circ = 60*PI cm diam = circum / PI = 60 radius = 30cm

OpenStudy (anonymous):

thank you wolf

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

When BA = 10 ft, find the area of the region that is NOT shaded. Round to the nearest whole number. 262 ft2 43 ft2 305 ft2 52 ft2 https://wcps6429-millbrookhs-ccl.gradpoint.com/Resource/3750901,780,111,0,0,0,0/Assets/testitemimages/geometry_b/area/mc114-1.jpg

OpenStudy (anonymous):

last one

OpenStudy (wolf1728):

Wow - just surfing the Internet just got back

OpenStudy (wolf1728):

BA=10 radius = 10 area of entire circle = PI*r² = 314.159 area of 60° sector = 60/360 * 314.159 = 52.35983 Area of shaded area (called a segment) = 9.0586 Area = Sector Area 52.35983 MINUS Segment Area 9.0586 Area = 43.301 To find segment area try this calculator: (yeah my website) http://www.1728.org/circsect.htm To figure out segment area you'll have to go here: http://www.1728.org/circpart.htm Hope that helps.